40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 23 15:27:19 2008, Alberto, Update, General, Abs length TRX_aplot_03_04_05_together.png
    Reply  Thu Oct 23 18:23:47 2008, Alberto, Update, General, Abs length TRX_plot_06_07_08_together.png
    Reply  Thu Oct 23 18:27:04 2008, Alberto, Update, General, Abs length HOM_Frequencies.m
Message ID: 1073     Entry time: Thu Oct 23 18:23:47 2008     In reply to: 1072
Author: Alberto 
Type: Update 
Category: General 
Subject: Abs length 

Quote:
Here are the measurements I've got yesterday. The plot shows the transmitted power after the X arm while sweeping the frequency of the beat between the two lasers. That frequency is changed by scanning the frequency of the local oscillator of the PLL (that is the Marconi).
The X arm cavity has been locked to the TEM00 of the main beam. I tilted ITMX in order to enhance the higher modes of the secondary beam with the purpose of making them beat with the main beam.
Three traces are shown in the plot correspondent to three different measurements in which I clipped the transmitted beam at the X end with a razor blade from up and from the side of the photodiode.
Both the beats of the TEM00 mode of the main laser with the TEM01 and TEM10 modes of the secondary laser are expected to be at 6.2763 MHz. The plot has a candidate peak at 6.325MHz but it does not appear on both the measurements with the blade. the peaks at 3.897MHz and 7.795MHz are the first and the second longitudinal modes of the X arm cavity.


Today I repeated the measurement and I'm attaching the resulting plot. Still, not clear and (and most of all) not nice.
It seems like tilting ITMX is introducing a lot of unwanted higher modes that don't let us to clearly identify TEM01 and TEM10.
I think I'm going to stop it to get back to technique in which the arm cavity is locked to the TEM01/10 of the main beam.
Attachment 1: TRX_plot_06_07_08_together.png  14 kB  | Hide | Hide all
TRX_plot_06_07_08_together.png
ELOG V3.1.3-